OpenStudy (babyslapmafro):

Please help me solve the given differential equation using the Cauchy-Euler equation: xy"+y'=0

OpenStudy (anonymous):

Assume \(y=x^r\) is a solution. \[y=x^r\\ y'=rx^{r-1}\\ y''=r(r-1)x^{r-2}\] Plug into the equation, \[x\left(r(r-1)x^{r-2}\right)+rx^{r-1}=0\] and solve for \(r\).